OpenStudy (anonymous):

Find the y-intercept and the x-intercept of the equation -25x+y=35

OpenStudy (anonymous):

25x+y=35 25(0)+y=35 y=35 (0,35) 25x+0=35 x=35/25 x=7/5 (7/5,0)